OpenStudy (anonymous):

Find 3x^2 - y^3 - y^3 - z if x = 3, y = -2, and z = -5. Numerical Answers Expected! I got 92, is that right?

OpenStudy (anonymous):

Pretty sure it's wrong though lol

OpenStudy (mokeira):

nope...not really

OpenStudy (anonymous):

I got 48

OpenStudy (mrnood):

I think it is wrong Just break it down... What is 3x^2 if x=3 What is y^3 if y= -2 don't forget you are subtracting y^3 in the equation....

OpenStudy (mokeira):

\[3x^2-y^3-y^3-z\] so, we substitute \[3(3^2)-(-2^3)-(-2^3)-(-5)=3(9)-(-8)-(-8)-(-5)\] \[27+8+8+5=48\]

OpenStudy (anonymous):

o3o welp I'm going to submit this lol

OpenStudy (anonymous):

and it was right! :p

OpenStudy (mrnood):

My reply was to your first answer of 92 - sorry for confusion. 48 is correct